# On the long-time limit of the mean curvature flow in closed manifolds

### v1

@@ 2.2. Ilmanen’s regularity theorem
Because the mean curvature flow is the gradient flow of area functional, the long-time limit is expected to be a critical point to the area functional. From geometric measure theory, the limit is known to be a stationary varifold, and without any other assumptions, the limit can have singularities.
In [30], Ilmanen used Simon’s sheeting theorem [41] to show that the tangent flow of a first-time singularity of a mean curvature flow of closed embedded surfaces in ${\mathbb{R}}^{3}$ must support on a smooth embedded shrinker, possibly with multiplicity. Ilmanen’s argument is very general and can be adapted to study the long-time limit of mean curvature flow. For the sake of clarity we precisely state the version that we will use in this paper:
Theorem 2.8 (Ilmanen’s regularity for limit surfaces).
Suppose $\{M_{i}\}_{i\in{\mathbb{Z}}_{+}}$ is a sequence of closed embedded surfaces in a closed manifold $(N^{3},g)$, satisfies the following assumptions:
@@
Then there exists a subsequence of $\{M_{i}\}_{i\in{\mathbb{Z}}_{+}}$, still denoted by $\{M_{i}\}$, that converges to an integral varifold $V$ in the sense of varifold, such that $V$ is supported on a finitely many disjoint closed embedded minimal surface $\Sigma_{1},\cdots,\Sigma_{k}$, with multiplicity $m_{1},\cdots,m_{k}\geq 1$. Moreover, for any $j=1,2,\cdots,k$, there exists finitely many points $p_{j,1},\cdots,p_{j,\ell_{j}}$ inside $\Sigma_{j}$, such that for any $r>0$ that is sufficiently small, for sufficiently large $i$, $M_{i}\backslash\bigcup_{l=1}^{\ell_{j}}B_{r}(p_{j,l})$ is the union of $m_{k}$ number of disjoint connected components, and each component is isotopic to and converges in the varifold sense with either multiplicity $1$ or $2$ to $\Sigma_{j}\backslash\bigcup_{l=1}^{\ell_{j}}B_{r}(p_{j,l})$ depending on if the limit is two sided or not.
@@
In this section we show that in the long term an almost regular flow either goes extinct in finite time or limits to a collection of smooth, but possibly unstable, minimal surfaces. Note that almost regular flows are not necessarily generic, although in the further applications we will often suppose our flow is – for posterities sake, we work in this more general class as much as possible. We finish with some properties of the minimal surfaces we find should the limit be nonempty. We start with the following total curvature bound:
@@
$\int_{M_{t}}|A|^{2}=\int_{M_{t}}H^{2}-4\pi\chi(M_{t})+C\text{Vol}(M_{t})$ (3.1)
Where the constant $C$, independent of $t$, depends on the curvature of $N$. By the property of almost regular flow [4], the genus of $M_{t}$ is bounded by a uniform constant $G$, and hence $\chi(M_{t})$ is also bounded by a constant. Of course $\text{Vol}(M_{t})$ is bounded in terms of $\text{Vol}(M)$. Since the time derivative of the area is given by $-\int_{M_{t}}H^{2}$ (for weak flows this is a consequence of Brakke’s inequality) we see measure of the set of times for which $\int H^{2}$ is greater than a constant $C_{1}$ is bounded by $C_{1}/\text{area}(M_{0})$. These two observations combined give what we want. ∎
@@
Of course, from general theory $\operatorname{supp}V$ is not necessarily a regular minimal surface. However, for surfaces in $3$–manifolds, we have more powerful tools to pick a convergence subsequence whose limit is regular, using Ilmanen’s regularity theory [30]. We adopt his argument here; Lemma 3.1 gives us the total curvature bounds we require in the application of Simon’s sheeting theorem, but there is a central feature of mean curvature flows in ${\mathbb{R}}^{3}$ that does not hold for mean curvature flows in a $3$–manifold which requires us to take some extra care. Namely Huisken’s monotonicity formula shows that a mean curvature flow $M_{t}$ of surfaces in ${\mathbb{R}}^{3}$ has a uniform area growth bound $\operatorname{Area}(M_{t}\cap B_{R}(x))\leq D\pi R^{2}$ for all $R>0$, $x\in{\mathbb{R}}^{3}$ and $t>0$, where $D$ only depends on $M_{0}$; see Proposition 3.3 in [44]. This area growth bound is crucial in applying Simon’s sheeting theorem in [30]; in particular see Theorem 10 and Corollary 11 of [30].
When $M_{t}$ is a mean curvature flow in a closed manifold $(N,g)$, by isometric embedding $(N,g)$ into ${\mathbb{R}}^{m}$, $M_{t}$ is a mean curvature flow with additional forces $\beta$, where $\beta$ is a tangent vector field on $N$, depending on the second fundamental form of $(N,g)$ in ${\mathbb{R}}^{m}$. Because of the forcing term, Huisken’s monotonicity formula can not directly provide uniform area growth bound for all future time $t$, but luckily a modified version of it can for short times which will suffice for our purposes.
To start, let us fix an isometric embedding of $(N,g)$ into ${\mathbb{R}}^{m}$; recalling that we suppose $N$ is compact and denote by $L<\infty$ an upper diameter bound for $N$ under this embedded. Then it happens that the forcing term $\beta$, $\beta(x)\in T_{x}N$, only depends on the second fundamental form of the isometric embedding of $(N,g)$. Hence $\|\beta\|_{L^{\infty}}\leq C$ for some fixed constant $C$. Considering
@@
which is the backward heat kernel in $2$-dimensions, the monotonicity formula of mean curvature flow with additional forces [51, 36] shows that for $0\leq t_{1}\leq t_{2}<t_{0}$,
[…]
We remark that the monotonicity formula also works for Brakke flows with additional forces, so this inequality holds even $M_{t}$ is not necessarily smooth for $t\in[t_{1},t_{2}]$. The following lemma shows that, while we may not get a uniform area growth bound for all $t>0$, we can show that for sufficiently large $t$, $M_{t}$ has a uniform area growth bound.
@@
[…]
@@
If the area of $M_{t}$ is sufficiently small the clearing out lemma [29, 12.2 and 12.5] shows that $M_{t}$ goes extinct (i.e. is supported on a set of zero measure) a short time later, so we may suppose its area is uniformly bounded below for all finite times. Considering a sequence $t_{i}\to\infty$ as in Lemma 3.2 above, by Allard compactness we may extract a converging subsequence which will be a stationary varifold $V$. By the almost all time regularity of the almost regular flows, we may suppose this set of times is regular as well. We also have the area growth estimate of $M_{t_{i}}$ from Lemma 3.4. With this in hand we may employ Simon’s sheeting theorem ala Ilmanen [30] along a sequence of smooth times $t_{i}\to\infty$ of $M_{t_{i}}$ to say that $V$ supports on a finite collection of smooth, embedded minimal surfaces $\{\Sigma_{1},\cdots,\Sigma_{k}\}$ and the convergence is $m_{j}$–sheeted away from a finite number of “bridge points” towards $\Sigma_{j}$. By the regularity of the limit, and the embeddedness of $M_{t_{i}}$, all the $\Sigma_{j}$’s are disjoint via the maximum principle. ∎
Higher multiplicity convergence can indeed occur, see [14], and in the case one of the $\Sigma_{i}$ is strictly stable should be a robust phenomenon. Also, it is easy to see the limiting set could be disconnected. Take for instance the initial data given by two disconnected area-minimizing closed embedded (say, strictly stable) minimal surfaces connected by a very thin neck. Then shortly after the flow is started the neck will quickly pinch and retract to each minimizing surface, and the limit of the flow will be the union of these two disconnected minimal surfaces.
It is interesting to ask to what extent the limit is unique. If the convergence has multiplicity $1$, and the manifold is analytic, then by the classic Łojasiewicz-Simon inequality, together with the Brakke/White regularity theory, $M_{t}$ indeed convergence to $\Sigma$ as $t\to\infty$, not only subsequentially. Without this machinery even in the most well-known case, the curve shortening flow on surfaces, this question is not completely understood. In fact in [21] Grayson conjectured that there exists an example of curve shortening flow whose limit geodesics are not unique.
It is not so hard though to see the limit will be unique if the limiting minimal surfaces are strictly stable, which we show next. To start the following lemma shows that if we carefully choose $t_{i}$, $M_{t_{i}}$ will stay in a neighborhood of the limit surfaces. In other words, those necks described in the example above indeed must disappear in a very short period of time.
@@
In the context of Corollary 1.1 given a sequence $s_{i}\to\infty$ along which the flow converges to a collection of minimal surfaces $\Sigma_{j}$ for any $\delta>0$ we can choose another sequence of $t_{i}\to\infty$ such that, when $t_{i}$ is sufficiently large, $M_{t_{i}}$ must lie inside the $\delta$-tubular neighborhood of $\cup_{j=1}^{k}\Sigma_{j}$.
@@
Denote by $\mathcal{N}_{\delta}(\Sigma)$ the (open) $\delta$-tubular neighborhood of a closed embedded surface $\Sigma$ in $N$, and consider a sequence of $s_{i}\to\infty$ as in Corollary 1.1 such that $M_{s_{i}}$ converges to $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ in the sense of varifolds. For any $x_{0}\in N\backslash(\bigcup_{j=1}^{k}\mathcal{N}_{\delta}(\Sigma_{j}))$, $\operatorname{Area}(M_{s_{i}}\cap B_{\delta}(x_{0}))\to 0$ because $B_{\delta}(x_{0})\cap(\bigcup_{j=1}^{k}\Sigma)=\emptyset$. In particular, this shows that for any $\varepsilon>0$, $x^{\prime}\in B_{\delta/2}(x_{0})$, and $\eta\in(1/2,3/4)$, the Gaussian weight
@@
Then by (3.3) again, we see that when $\varepsilon$ is chosen sufficiently small, the Gaussian density of $x^{\prime}\in B_{\delta/2}(x_{0})$ of $M_{t}$ for $t\in(s_{i}+\delta/2,s_{i}+3\delta/4)$ is strictly less than $1$, hence must be 00. In other words, for $t\in(s_{i}+\delta/2,s_{i}+3\delta/4)$, $M_{t}$ is disjoint from $B_{\delta/2}(x_{0})$ for sufficiently large $s_{i}$. Because $N\backslash\bigcup_{j=1}^{k}\mathcal{N}_{\delta}(\Sigma_{j})$ is compact, we can cover it by finitely many $B_{\delta/2}(x_{0})$’s. This allows us to pick a sequence of $t_{i}\to\infty$ such that $M_{t_{i}}\subset\bigcup_{j=1}^{k}\mathcal{N}_{\delta}(\Sigma_{j})$ that have bounded total curvature and total mean curvature tending to zero. The argument of Corollary 1.1 may then be rerun on this sequence to extract a subsequence which converges to a union of smoothly embedded minimal surfaces $\Sigma_{j}^{\prime}$. Using these minimal surfaces the conclusion follows by the triangle inequality. ∎
We remark that in the above proof conceivably $\Sigma_{i}\neq\Sigma_{i}^{\prime}$, and in particular the further chosen subsequence $M_{t_{j}}$ may not converge to the same collection of closed embedded minimal surfaces $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$. But their limit must be very close to them, in the sense that the limit lies in a tubular neighborhood of these minimal surfaces. Now, recall that a stable minimal surface $\Sigma$ is strictly stable if the linearized operator $L_{\Sigma}$ has only positive eigenvalues. Using a barrier argument and the lemma above then we observe we may show that if the limit stable minimal surfaces are strictly stable, then the limit is unique, even if the convergence is possibly in high multiplicity and in a merely smooth (but not analytic) background manifold $N$:
@@
Suppose $M_{t}$ is an almost regular flow such that there exists $t_{i}\to\infty$ such that $M_{t_{i}}\to m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ as in Theorem 1.1, and each $\Sigma_{j}$ is strictly stable. Then the limit is unique: namely for any sequence of $s_{i}\to\infty$ so that $M_{s_{i}}$ converges as varifolds the limit must be $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$.
@@
For any $\delta>0$, we can apply Lemma 3.5 to claim that there is another sequence, which we still label $t_{i}$, such that $M_{t_{i}}$ lies inside the union of $\delta$-tubular neighborhood of the $\Sigma_{j}$’s when $t_{i}$ is sufficiently large.
@@
To proceed, we use the neighborhood foliation structure of strictly stable minimal surfaces: because each $\Sigma_{j}$ is strictly stable, if it is $2$-sided, then there exists a tubular neighborhood $\mathcal{N}^{j}$ of $\Sigma_{j}$ such that $\mathcal{N}^{j}$ is foliated by surfaces $\Sigma^{t}$ for $t\in(-s,s)$, such that when $t\in(-s,0)\cup(0,s)$, $\Sigma^{t}$ has positive (inward pointing) mean curvature. For $1$-sided $\Sigma_{j}$, there is a similar structure of the neighborhood. For a proof of this fact, see [42, Lemma A.1] for $2$-sided case, and [42, Section 6] for $1$-sided case.
@@
Using this fact, we can find a tubular neighborhood $\mathcal{N}^{j}$ of $\Sigma_{j}$ such that $\partial\mathcal{N}^{j}$ has strictly positive mean curvature. In particular, if at the beginning we choose $\delta>0$ sufficiently small so that the $\delta$-tubular neighborhood is contained in $\mathcal{N}^{j}$, then $M_{t}$ will lie within $\bigcup_{j=1}^{k}\mathcal{N}^{j}$ for $t\geq t_{i}$. Hence we see any limit of $M_{t}$ has to be supported in $\bigcup_{j=1}^{k}\mathcal{N}^{j}$. If $\delta>0$ is sufficiently small we may apply Brakke regularity to see the flows of boundary components of $\bigcup_{j=1}^{k}\mathcal{N}^{j}$ are smooth for all $t>0$ and asymptote back to $\bigcup\Sigma_{j}$, so by the comparison principle any converging limit of $M_{t}$ has to be supported on $\bigcup_{j=1}^{k}\Sigma_{j}$. The question of convergence in varifold topology to this entire set is more subtle though, because in the above we may only apply Ilmanen’s analysis to special sequence of times $s_{i}\to\infty$.
@@
To deal with this, first we point out that for any sequence $s_{i}\to\infty$ the argument from the proof of Theorem 1.1 does apply we get the limit must be supported precisely on $\bigcup_{j=1}^{k}\Sigma_{j}$. Moreover by the monotonicity of area along the mean curvature flow the multiplicities may not vary either, so that for any sequence $t_{i}\to\infty$ for which $M_{t_{i}}$ converges "well" as in the proof of Theorem 1.1, $M_{t_{i}}\to m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ as $t_{i}\to\infty$. Now, for a given $1\leq j\leq k$ choose a small $\epsilon>0$ let $u_{j,\epsilon,D}$ be a smooth function which is supported on the $\epsilon$ tubular neighborhood of a domain $D\subset\Sigma_{j}$. Because almost regular flows are smooth for almost all times and continuous in the weak topology in time (see Lemma 2.7 in [4]) Lemmas 3.1, 3.2 gives us that for any $\delta>0$ we can approximate any sequence $s_{i}\to\infty$ by another sequence $s_{i}^{\prime}\to\infty$ so that $\int_{M_{s_{i}}}u_{j,\epsilon,D}-\int_{M_{s_{i}^{\prime}}}u_{j,\epsilon,D}<\delta$ and that Ilmanen’s analysis applies to the sequence $s_{i}^{\prime}$. Because from the above any converging limit of $M_{t}$ must be contained in $\bigcup\Sigma_{j}$ if the sequence of $M_{s_{i}}$ converges as varifolds its support must be contained in $\bigcup_{j=1}^{k}\Sigma_{j}$, and so by ranging over $D,j$ and letting $\epsilon,\delta\to 0$ we thus get that if the sequence $M_{s_{i}}$ converges as varifolds the limit must be precisely $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ as claimed. ∎
@@
Indeed if one of the minimal surfaces $\Sigma_{i}$ is merely stable but not strictly stable, the above contracting tubular neighborhood may not exist. An example of such a minimal surface can be found in [42, Appendix B]. To wrap up this section we summarize some topological properties of the limit minimal surfaces in the case $M_{t}$ can be approximated by generic flows:
@@
Suppose that $M_{t}$ is an inner or outermost flow of $M$ or is a generic flow whose long term limit is nonempty, and let $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ be limiting minimal surfaces as in Theorem 1.1, and that $N$ is oriented implying 2–sidedness is equivalent to orientability. Then the following hold:
@@
    $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ in Corollary 1.1 is homologous to $M$ in $H_{2}(N,{\mathbb{Z}}_{2})$.
@@
    If $\Sigma_{j}$ is $1$–sided, then $m_{j}$ is even.

### v2

@@ 2.2. Ilmanen’s regularity theorem
Because the mean curvature flow is the gradient flow of the area functional, the long-time limit is expected to be a critical point of the area functional. From geometric measure theory, the limit is known to be a stationary varifold, and without any other assumptions, the limit can have singularities.
In [Ilmanen95_Sing2D, Section 4, proof of Theorem 1.1], Ilmanen used Simon’s sheeting theorem [Simon94_Willmore] to show that the tangent flow of a first-time singularity of a mean curvature flow of closed embedded surfaces in ${\mathbb{R}}^{3}$ must support on a smooth embedded shrinker, possibly with multiplicity. Ilmanen’s argument is very general and can be adapted to study the long-time limit of mean curvature flow. For the sake of clarity, we precisely state the version that we will use in this paper:
Theorem 2.8 (Ilmanen’s regularity for limit surfaces, [Ilmanen95_Sing2D], Section 4, proof of Theorem 1.1; see also [BamlerKleiner23_Multiplicity1], Lemma 2.13).
Suppose $\{M_{i}\}_{i\in{\mathbb{Z}}_{+}}$ is a sequence of closed embedded surfaces in a closed manifold $(N^{3},g)$, satisfying the following assumptions:
@@
[…]
@@
In this section, we show that in the long term, an almost regular flow either goes extinct in finite time or limits to a collection of smooth, but possibly unstable, minimal surfaces. Note that almost regular flows are not necessarily generic, although in further applications we will often suppose our flow is – for posterities sake, we work in this more general class as much as possible. We finish with some properties of the minimal surfaces we find, should the limit be nonempty.
Although as stated, the following lemma is not required to invoke Theorem 2.8 below, the reader may recall that an important step in its proof in [Ilmanen95_Sing2D] is to show that the surfaces involved have finite total curvature (see the proof of Theorem 1.1 below). With this in mind, we start with the following total curvature bound for the sake of exposition:
@@
$\int_{M_{t}}|A|^{2}=\int_{M_{t}}H^{2}-4\pi\chi(M_{t})+C\text{Vol}(M_{t}).$ (3.1)
Where the constant $C$, independent of $t$, depends on the curvature of $N$. By the property of almost regular flow [BamlerKleiner23_Multiplicity1], the genus of $M_{t}$ is bounded by a uniform constant $G$, and hence $\chi(M_{t})$ is also bounded by a constant. Of course $\text{Vol}(M_{t})$ is bounded in terms of $\text{Vol}(M)$. Since the time derivative of the area is given by $\int_{M_{t}}H^{2}$ (for weak flows, this is a consequence of Brakke’s inequality), we see measure of the set of times for which $\int H^{2}$ is greater than a constant $C_{1}$ is bounded by $C_{1}/\text{area}(M_{0})$. These two observations combined give what we want. ∎
@@
Of course, from general theory, $\operatorname{supp}V$ is not necessarily a regular minimal surface. However, for surfaces in 33–manifolds, we have more powerful tools to pick a convergence subsequence whose limit is regular, using Ilmanen’s regularity theory, Theorem 2.8. We adopt his argument here; Lemma 3.1 gives us the total curvature bounds we require in the application of Simon’s sheeting theorem, but there is a central feature of mean curvature flows in ${\mathbb{R}}^{3}$ that does not hold for mean curvature flows in a 33–manifold, which requires us to take some extra care. Namely Huisken’s monotonicity formula shows that a mean curvature flow $M_{t}$ of surfaces in ${\mathbb{R}}^{3}$ has a uniform area growth bound $\operatorname{Area}(M_{t}\cap B_{R}(x))\leq D\pi R^{2}$ for all $R>0$, $x\in{\mathbb{R}}^{3}$ and $t>0$, where $D$ only depends on $M_{0}$; see Proposition 3.3 in [sun-generic-multi-1]. This area growth bound is crucial in applying Simon’s sheeting theorem in [Ilmanen95_Sing2D]; in particular, see Theorem 10 and Corollary 11 of [Ilmanen95_Sing2D].
When $M_{t}$ is a mean curvature flow in a closed manifold $(N,g)$, by isometric embedding $(N,g)$ into ${\mathbb{R}}^{m}$, $M_{t}$ is a mean curvature flow with additional forces $\beta$, where $\beta$ is a tangent vector field on $N$, depending on the second fundamental form of $(N,g)$ in ${\mathbb{R}}^{m}$. Because of the forcing term, Huisken’s monotonicity formula can not directly provide uniform area growth bound for all future time $t$, but luckily, a modified version of it can for short times, which will suffice for our purposes.
To start, let us fix an isometric embedding of $(N,g)$ into ${\mathbb{R}}^{m}$; recalling that we suppose $N$ is compact and denote by $L<\infty$ an upper diameter bound for $N$ under this embedding. Then it happens that the forcing term $\beta$, $\beta(x)\in T_{x}N$, only depends on the second fundamental form of the isometric embedding of $(N,g)$. Hence $\|\beta\|_{L^{\infty}}\leq C$ for some fixed constant $C$. Considering
@@
which is the backward heat kernel in $2$-dimensions, the monotonicity formula of mean curvature flow with additional forces [White97_Stratif, AM_CMgenericambient] shows that for $0\leq t_{1}\leq t_{2}<t_{0}$,
$\int_{M_{t_{2}}}\rho_{x_{0},t_{0}}(x,t_{2})dx\leq e^{C(t_{2}-t_{1})}\int_{M_{t_{1}}}\rho_{x_{0},t_{0}}(x,t_{1})dx.$ (3.3)
We remark that the monotonicity formula also works for Brakke flows with additional forces, so this inequality holds even if $M_{t}$ is not necessarily smooth for $t\in[t_{1},t_{2}]$. The following lemma shows that, while we may not get a uniform area growth bound for all $t>0$, we can show that for sufficiently large $t$, $M_{t}$ has a uniform area growth bound.
@@
$\begin{split}\int_{M_{t-\mathbf{t}}}\rho_{x_{0},t+R^{2}}(x_{0},t-\mathbf{t})=&\int_{M_{t-\mathbf{t}}}(4\pi(R^{2}+\mathbf{t}))^{-1}e^{-\frac{|x-x_{0}|^{2}}{4(R^{2}+\mathbf{t})}}dx\\ \leq&(4\pi(R^{2}+\mathbf{t}))^{-1}\operatorname{Area}(M_{t-\mathbf{t}})\leq(4\pi\mathbf{t})^{-1}\operatorname{Area}(M_{0}).\end{split}$
@@
If the area of $M_{t}$ is sufficiently small, the clearing out lemma [Ilmanen94_EllipReg, 12.2 and 12.5] shows that $M_{t}$ goes extinct (i.e. is supported on a set of zero measure) a short time later. Therefore, we may suppose its area is uniformly bounded below for all finite times. Considering a sequence $t_{i}\to\infty$ as in Lemma 3.2 above, by Allard compactness we may extract a converging subsequence which will be a stationary varifold $V$. By the almost all time regularity of the almost regular flows, we may suppose this set of times is regular as well. We also have the area growth estimate of $M_{t_{i}}$ from Lemma 3.4, so that we may apply Theorem 2.8. For a brief overview, with the facts above in hand we may employ Simon’s sheeting theorem ala Ilmanen [Ilmanen95_Sing2D, Theorem 10 and Corollary 11] along a sequence of smooth times $t_{i}\to\infty$ of $M_{t_{i}}$ to say that $V$ supports on a finite collection of smooth, embedded minimal surfaces $\{\Sigma_{1},\cdots,\Sigma_{k}\}$ and the convergence is $m_{j}$–sheeted away from a finite number of "bridge points" towards $\Sigma_{j}$. By the regularity of the limit, and the embeddedness of $M_{t_{i}}$, all the $\Sigma_{j}$’s are disjoint via the maximum principle. ∎
Higher multiplicity convergence can indeed occur, see [ChenSun24_Multi2inManifold], and in the case one of the $\Sigma_{i}$ is strictly stable should be a robust phenomenon. Also, it is easy to see that the limiting set could be disconnected. Take, for instance, the initial data given by two disconnected area-minimizing closed embedded (say, strictly stable) minimal surfaces connected by a very thin neck. Then shortly after the flow is started, the neck will quickly pinch and retract to each minimizing surface, and the limit of the flow will be the union of these two disconnected minimal surfaces.
It is interesting to ask to what extent the limit is unique. If the convergence has multiplicity $1$, and the manifold is analytic, then by the classic Łojasiewicz-Simon inequality, together with the Brakke/White regularity theory, $M_{t}$ indeed convergence to $\Sigma$ smoothly as $t\to\infty$, not only subsequentially. Without this machinery, even in the most well-known case, the curve shortening flow on surfaces, this question is not completely understood. In fact, in [Grayson89_CSF] Grayson conjectured that there exists an example of curve shortening flow whose limit geodesics are not unique.
It is not so hard, though, to see that the limit will be unique if the limiting minimal surfaces are strictly stable, which we show next. To start, the following lemma shows that if we carefully choose $t_{i}$, $M_{t_{i}}$ will stay in a neighborhood of the limit surfaces. In other words, those necks described in the example above indeed must disappear in a very short period of time.
@@
In the context of Theorem 1.1 given a sequence $s_{i}\to\infty$ along which the flow converges to a collection of minimal surfaces $\Sigma_{j}$ for any $\delta>0$ we can choose another sequence of $t_{i}\to\infty$ such that, when $t_{i}$ is sufficiently large, $M_{t_{i}}$ must lie inside the $\delta$-tubular neighborhood of $\cup_{j=1}^{k}\Sigma_{j}$.
@@
Denote by $\mathcal{N}_{\delta}(\Sigma)$ the (open) $\delta$-tubular neighborhood of a closed embedded surface $\Sigma$ in $N$, and consider a sequence of $s_{i}\to\infty$ as in Theorem 1.1 such that $M_{s_{i}}$ converges to $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ in the sense of varifolds. For any $x_{0}\in N\backslash(\bigcup_{j=1}^{k}\mathcal{N}_{\delta}(\Sigma_{j}))$, $\operatorname{Area}(M_{s_{i}}\cap B_{\delta}(x_{0}))\to 0$ because $B_{\delta}(x_{0})\cap(\bigcup_{j=1}^{k}\Sigma)=\emptyset$. In particular, this shows that for any $\varepsilon>0$, $x^{\prime}\in B_{\delta/2}(x_{0})$, and $\eta\in(1/2,3/4)$, the Gaussian weight
$\int_{M_{s_{i}}}\rho_{x^{\prime},s_{i}+(\eta\delta)^{2}}(x,s_{i})dx<\varepsilon,\text{when $s_{i}$ is sufficiently large.}$
@@
We remark that in the above proof, conceivably $\Sigma_{i}\neq\Sigma_{i}^{\prime}$, and in particular, the further chosen subsequence $M_{t_{j}}$ may not converge to the same collection of closed embedded minimal surfaces $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$. But their limit must be very close to them, in the sense that the limit lies in a tubular neighborhood of these minimal surfaces. Now, recall that a stable minimal surface $\Sigma$ is strictly stable if the linearized operator $L_{\Sigma}$ has only positive eigenvalues. Using a barrier argument and the lemma above, we observe we may show that if the limit stable minimal surfaces are strictly stable, then the limit is unique, even if the convergence is possibly in high multiplicity and in a merely smooth (but not necessarily analytic) background manifold $N$:
@@
Suppose $M_{t}$ is an almost regular flow such that there exists $t_{i}\to\infty$ such that $M_{t_{i}}\to m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ as in Theorem 1.1, and each $\Sigma_{j}$ is strictly stable. Then the limit is unique: namely, for any sequence of $s_{i}\to\infty$ so that $M_{s_{i}}$ converges as varifolds, the limit must be $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$.
@@
For any $\delta>0$, we can apply Lemma 3.5 to claim that there is another sequence, which we still label $t_{i}$, such that $M_{t_{i}}$ lies inside the union of the $\delta$-tubular neighborhood of the $\Sigma_{j}$’s when $t_{i}$ is sufficiently large.
@@
To proceed, we use the neighborhood foliation structure of strictly stable minimal surfaces: because each $\Sigma_{j}$ is strictly stable, if it is $2$-sided, then there exists a tubular neighborhood $\mathcal{N}^{j}$ of $\Sigma_{j}$ such that $\mathcal{N}^{j}$ is foliated by surfaces $\Sigma^{t}$ for $t\in(-s,s)$, such that when $t\in(-s,0)\cup(0,s)$, $\Sigma^{t}$ has positive (inward pointing) mean curvature. For $1$-sided $\Sigma_{j}$, there is a similar structure of the neighborhood. For a proof of this fact, see [StevensSun24_LargeArea, Lemma A.1] for $2$-sided case, and [StevensSun24_LargeArea, Section 6] for $1$-sided case.
@@
Using this fact, we can find a tubular neighborhood $\mathcal{N}^{j}$ of $\Sigma_{j}$ such that $\partial\mathcal{N}^{j}$ has strictly positive mean curvature. In particular, if at the beginning we choose $\delta>0$ sufficiently small so that the $\delta$-tubular neighborhood is contained in $\mathcal{N}^{j}$, then $M_{t}$ will lie within $\bigcup_{j=1}^{k}\mathcal{N}^{j}$ for $t\geq t_{i}$. Hence, we see any limit of $M_{t}$ has to be supported in $\bigcup_{j=1}^{k}\mathcal{N}^{j}$. If $\delta>0$ is sufficiently small we may apply Brakke regularity to see the flows of boundary components of $\bigcup_{j=1}^{k}\mathcal{N}^{j}$ are smooth for all $t>0$ and asymptote back to $\bigcup\Sigma_{j}$, so by the comparison principle any converging limit of $M_{t}$ has to be supported on $\bigcup_{j=1}^{k}\Sigma_{j}$. The question of convergence in varifold topology to this entire set is more subtle, though, because in the above we may only apply Ilmanen’s analysis to a special sequence of times $s_{i}\to\infty$.
@@
To deal with this, first we point out that for any sequence $s_{i}\to\infty$, the argument from the proof of Theorem 1.1 does apply, we get the limit must be supported precisely on $\bigcup_{j=1}^{k}\Sigma_{j}$. Moreover, by the monotonicity of area along the mean curvature flow, the multiplicities may not vary either, so that for any sequence $t_{i}\to\infty$ for which $M_{t_{i}}$ converges "well" as in the proof of Theorem 1.1, $M_{t_{i}}\to m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ as $t_{i}\to\infty$. Now, for a given $1\leq j\leq k,$ choose a small $\epsilon>0$, let $u_{j,\epsilon,D}$ be a smooth function which is supported on the $\epsilon$ tubular neighborhood of a domain $D\subset\Sigma_{j}$. Because almost regular flows are smooth for almost all times and continuous in the weak topology in time (see Lemma 2.7 in [BamlerKleiner23_Multiplicity1]) Lemmas 3.1, 3.2 give us that for any $\delta>0$ we can approximate any sequence $s_{i}\to\infty$ by another sequence $s_{i}^{\prime}\to\infty$ so that $|\int_{M_{s_{i}}}u_{j,\epsilon,D}-\int_{M_{s_{i}^{\prime}}}u_{j,\epsilon,D}|<\delta$ and that Ilmanen’s analysis applies to the sequence $s_{i}^{\prime}$. Because from the above any converging limit of $M_{t}$ must be contained in $\bigcup\Sigma_{j}$ if the sequence of $M_{s_{i}}$ converges as varifolds its support must be contained in $\bigcup_{j=1}^{k}\Sigma_{j}$, and so by ranging over $D,j$ and letting $\epsilon,\delta\to 0$ we thus get that if the sequence $M_{s_{i}}$ converges as varifolds the limit must be precisely $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ as claimed. ∎
@@
Indeed, if one of the minimal surfaces $\Sigma_{i}$ is merely stable but not strictly stable, the above contracting tubular neighborhood may not exist. An example of such a minimal surface can be found in [StevensSun24_LargeArea, Appendix B]. To wrap up this section, we summarize some topological properties of the limit minimal surfaces in the case $M_{t}$ can be approximated by generic flows:
@@
Suppose that $M_{t}$ is an inner or outermost flow of $M$ or is a generic flow whose long term limit is nonempty, and let $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ be limiting minimal surfaces as in Theorem 1.1, and that $N$ is oriented, implying 2–sidedness is equivalent to orientability. Then the following hold:
@@
    $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ in Theorem 1.1 is homologous to $M$ in $H_{2}(N,{\mathbb{Z}}_{2})$.
@@
    If $\Sigma_{j}$ is 11–sided, then $m_{j}$ is even.